The initial speed of a bullet fired from a rifle is 630 m/s. The rifle is fired at the centre of a target 700 m away at the same level as the target. How far above the centre of the target ?

Grade:12th pass

1 Answers

Rahul Pattankar
7 years ago
We know that y=(xtan theta)-(gx^2)/2u^2cos^2theta.....(i) T=distance in x direction/speed in x direction=>T=700/630=>T=1.11Now according to trajectory of projectile,x=(u cos theta) TPutting x in (i)y=(u cos theta) T tan theta - (g u^2 cos^2 theta T^2)/2u^2 cos^2 thetaAs projectile is projected from a height with angle zero with x direction theta=0=>y=gT^2/2=>y=10×1.11×1.11/2=>y=5×1.11×1.11=>y=6.1m